A Bibliography of Publications about the REXX Scripting Language

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

   E-mail: beebe@math.utah.edu, beebe@acm.org,
              beebe@computer.org (Internet)
  WWW URL: http://www.math.utah.edu/~beebe/

06 May 2022 Version 1.19

References

IBM:1983:SPIa

[1] IBM Corporation, Endicott, NY, USA. System Product: Interpreter Reference, 1983.

IBM:1983:SPIb

[2] IBM Corporation, Endicott, NY, USA. System Product: Interpreter User’s Guide, 1983.

Brodock:1984:SPI

[3] G. Brodock and C. Morse. The system product interpreter (REXX) examples and techniques. Technical Bulletin GG22-9361-00, IBM Corporation, San Jose, CA, USA, 1984. viii + 62 pp.

Cowlishaw:1984:DRL

[4] Mike F. Cowlishaw. The design of the REXX language. IBM Systems Journal, 23(4):326–335, 1984. CODEN IBMSA7. ISSN

0018-8670. URL http://www.research.ibm.com/journal/sj/234/

ibmsj2304D.pdf. First published as IBM Hursley Technical Report TR12.223, October 1983.

Cowlishaw:1985:PR

[5] M. F. Cowlishaw. Personal REXX. Mansfield Software Group, Storrs, CT, USA, 1985. Computer diskette.

Cowlishaw:1985:RLP

[6] M. F. Cowlishaw. The REXX language: a practical approach to programming. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1985. ISBN 0-13-780735-X (paperback). xi + 176 pp. LCCN QA76.73.R24 C69 1985.

OHara:1985:MPU

[7] Robert P. O’Hara and David R. (David Roos) Gomberg. Modern Programming Using REXX. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1985. ISBN 0-13-597311-2. xii + 194 pp. LCCN QA76.73.R24 O33 1985.

Cowlishaw:1986:PR

[8] M. F. Cowlishaw. Personal REXX. Mansfield Software Group, Storrs, CT, USA, 1986. Computer diskette.

Cowlishaw:1987:DRL

[9] M. Cowlishaw. The design of the REXX language. ACM SIGPLAN Notices, 22(2):26–35, February 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cowlishaw:1987:RGN

[10] Mike F. Cowlishaw, Yoichi Muraoka, and Koji Shibano. Rekkusu gengo nyumon: [puroguraminguhe no jissenteki apurochi], volume 17 of Sofutowea kogaku raiburari. Kindai Kagakusha, Tokyo, Japan, 1987. ISBN 4-7649-0136-6. 172 pp. Japanese translation of [6].

Hawes:1987:AUR

[11] William S. Hawes. ARexx user’s reference manual, version 1.0: the REXX language for the Amiga. ????, ????, 1987. 155 pp. LCCN QA76.73.A275.

IBM:1987:GRG

[12] IBM Corporation, San Jose, CA, USA. GDDM: REXX guide, 1987. vii + 78 pp. Licensed Program. Release 1. SC33-0478-0. Program Number GDDM-REXX 5664-336. File No. S370/4300/CMS-34.

IBM:1987:WSP

[13] IBM Corporation, San Jose, CA, USA. Writing simple programs with REXX: VM/Integrated System, release 5, 1987. xiv + 137 pp. SC24-5357-00. File no. 5370/4300-40.

Cowlishaw:1988:PR

[14] Mike F. Cowlishaw. Die Programmiersprache REXX (German) [The REXX Programming Language]. Carl Hanser, München, Germany, 1988. ISBN 3-446-15195-8 ; 0-13-780784-8. xii + 180 pp. LCCN ????

IBM:1988:TEVa

[15] IBM Corporation, San Jose, CA, USA. TSO extensions Version 2: REXX reference, December 1988. various pp. SC28-1883-0. File number S370-39. This edition applies to the TSO Extensions (TSO/E) Version 2 Licensed Program, Program Number 5685-025, and to all subsequent releases until otherwise indicated in new editions or Technical Newsletters.

IBM:1988:TEVb

[16] IBM Corporation, San Jose, CA, USA. TSO extensions Version 2: REXX user’s guide, 1988. various pp. SC28-1882-0. File number S370-39. Program number 5685-025. This edition applies to the TSO Extensions (TSO/E) Version 2 Licensed Program, Program Number 5685-025, and to all subsequent releases until otherwise indicated in new editions or Technical Newsletters.

Nunn:1988:RPC

[17] Debbie Nunn. The REXX project: a case study of automatic hypertext construction. Technical report TAMU 88-021, Texas A & M University, Computer Science Department, College Station, TX, USA, April 1988. vi + 72 pp.

OHara:1988:MPU

[18] Robert P. O’Hara and David Roos Gomberg. Modern Programming Using REXX. Prentice-Hall, Upper Saddle River, NJ 07458, USA, revised edition, 1988. ISBN 0-13-597329-5. xiii + 218 pp. LCCN QA76.73.R24 O33 1988. US$21.00.

Weintraub:1988:ACS

[19] D. M. Weintraub. APL2 and the CMS system: Exploiting the APL2/ REXX connection. In Dickey and Shaw [127], pages 337–345. ISBN 0-89791-253-5. ISSN 0163-6006. LCCN ???? URL http://www.acm.org/

pubs/citations/proceedings/apl/55626/p337-weintraub/. ACM or-

der number: 554880.

Hoernes:1989:RT

[20] Gerhard E. Hoernes. REXX on TSO/E. IBM Systems Journal, 28(2): 274–293, 1989. CODEN IBMSA7. ISSN 0018-8670.

IBM:1989:TEV

[21] IBM Corporation, San Jose, CA, USA. TSO extensions Version 2: REXX reference, second edition, December 1989. various pp. SC28-1883-1. File number S370-39.

Anonymous:1990:VDX

[22] Anonymous. V/DEVELOPER XMENU/REXX interface user’s guide and reference: version 2 release 3. VM Systems Group Inc., Vienna, VA, USA, 1990. viii + 190 pp. LCCN QA76.9.D6.

Cowlishaw:1990:RLP

[23] Michael F. Cowlishaw. The REXX language: a practical approach to programming. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1990. ISBN 0-13-780651-5. xii + 203 pp. LCCN QA76.73.R24 C69

1990. URL http://vig.prenhall.com/catalog/academic/product/0,
1144,0137806515,00.html.

Daney:1990:RCR

[24] C. Daney. Rexx in charge: Rexx now can control and coordinate all aspects of the OS/2 environment. Byte Magazine, 15(8):245–246, 248, 250, 252–253, August 1990. CODEN BYTEDJ. ISSN 0360-5280.

Gargiulo:1990:RTE

[25] Gabriel F. Gargiulo. REXX in the TSO environment. QED IBM Mainframe series. QED Information Sciences, Inc., P. O. Box 82-181, Wellesley, MA 02181, USA, 1990. ISBN 0-89435-354-3. xxi + 318 pp. LCCN QA76.73.R24 G37 1990. US$39.95.

IBM:1990:TEVa

[26] IBM Corporation, San Jose, CA, USA. TSO extensions version 2: procedures language MVS/REXX user’s guide, second edition, October 1990. SC28-1882-1. File number: S370-39. Program number: 5685-025.

IBM:1990:TEVb

[27] IBM Corporation, San Jose, CA, USA. TSO extensions version 2: procedures language MVS/REXX reference, October 1990. various pp. SC28-1883-2. File number: S370-39.

Rudd:1990:PUR

[28] Anthony S. Rudd. Practical usage of REXX. Ellis Horwood series in computers and their applications. Ellis Horwood, New York, NY, USA, 1990. ISBN 0-13-682790-X. 294 pp. LCCN QA76.73.R24 R83 1990. US$50.95.

WG:1990:URR

[29] The Workstation Group, 525 Capital Drive, Lake Zurich, IL 60047, USA. uni-REXX Reference Manual, 1990. x + 404 pp. URL http:/

/www.wrkgrp.com/support/manuals/rxref297.pdf. Document number

TXV2-9.

Yaghi:1990:IRE

[30] H. M. Yaghi and J. E. Cross. Introducing REXX into the engineering curriculum. In IEEE [129], pages 825–827 (vol. 3). ISBN ???? LCCN ???? Three volumes.

Abraham:1991:TSS

[31] D. G. Abraham, G. M. Dolan, G. P. Double, and J. V. Stevens. Transaction Security System. IBM Systems Journal, 30(2):206–229, 1991. CO-DEN IBMSA7. ISSN 0018-8670. See erratum [32].

Anonymous:1991:ESS

[32] Anonymous. Erratum: “Transaction Security System”. IBM Systems Journal, 30(4):598, 1991. CODEN IBMSA7. ISSN 0018-8670. See [31].

Cowlishaw:1991:PRU

[33] M. F. Cowlishaw. Personal REXX user’s guide (PC-DOS and OS/2 REXX) version 3.0. Quercus Systems, 1991.

Giguere:1991:RJW

[34] Eric Giguère. Rexx: Not just a wonder dog. Computer Language Magazine, 8(3):45–??, March 1991. CODEN COMLEF. ISSN 0749-2839.

IBM:1991:TEV

[35] IBM. TSO extensions version 2: procedures language MVS/REXX reference. IBM Corporation, San Jose, CA, USA, 1991. xxii + 540 pp. SC28-1883-4. File number: S370/S390-39.

IBM:1991:TEVa

[36] IBM Corporation, San Jose, CA, USA. TSO extensions version 2: system diagnosis: REXX and CLIST, third edition, March 1991. various pp. LY28-1887-2. Program number: 5685-025.

IBM:1991:TEVb

[37] IBM Corporation, San Jose, CA, USA. TSO extensions version 2: procedures language MVS/REXX user’s guide, third edition, 1991. xiv + 248 pp. SC28-1882-2. File number: S370/S390-39. Program number: 5685-025.

IBM:1991:TEVc

[38] IBM Corporation, San Jose, CA, USA. TSO extensions version 2: procedures language MVS/REXX reference, fourth edition, March 1991. various pp. SC28-1883-3. File number: S370/S390–39. Program number: 5685-025.

Pinter:1991:PCR

[39] Ron Y. Pinter, Pnina Vortman, and Zvi Weiss. Partial compilation of REXX. IBM Systems Journal, 30(3):312–321, 1991. CODEN IBMSA7. ISSN 0018-8670.

Varian:1991:VVC

[40] Melinda Varian. VM and the VM community: Past, present, and future. Technical report, Office of Computing and Information Technology, Princeton University, Princeton, NJ 08544, USA, April 1991. 168 pp. URL

http://www.leeandmelindavarian.com/Melinda/neuvm.pdf. Original

presented at Australasian SHARE/GUIDE in Melbourne, Victoria, Australia in 1989. This is a detailed history of the development of virtual machine technology on IBM System/360 and later mainframes, and of the opposition by much of IBM to that technology until it was demonstrated that their performance could equal, or even exceed, that of an operating system running on bare hardware, and also allow a single physical host to support multiple operating systems, and software development, simultaneously. There are also several comments about the development of the REXX language, and about the influence of Unix on IBM’s software development.

Watts:1991:PRM

[41] Keith Watts. Portable/REXX for MS/DOS. Kilowatt Software, ????, 1991. ISBN ???? ???? pp. LCCN ???? Guide, Reference manual, Examples Reference, Reference Summary, and Learning to Program with Portable/REXX.

Watts:1991:RW

[42] Keith Watts. REXX/Windows. Kilowatt Software, ????, 1991. ISBN ???? ???? pp. LCCN ???? Product Guide and Reference.

Zamara:1991:UAA

[43] Chris Zamara and Nick Sullivan. Using ARexx on the Amiga. Abacus Software, Inc., ????, 1991. ISBN 1-55755-114-6. 424 pp. LCCN ???? Includes diskette.

Callaway:1992:AC

[44] Merrill Callaway. The ARexx Cookbook. Whitestone, ????, 1992. ISBN 0-9632773-0-8 (book), 0-9632773-1-6 (diskette). 221 pp. LCCN ????

Cowlishaw:1992:PRU

[45] M. F. Cowlishaw. Personal REXX user’s guide, version 3.0 — OS/2 supplement. Quercus Systems, 1992.

Daney:1992:PR

[46] Charles Daney. Programming in REXX. J. Ranade IBM series. McGraw-Hill, New York, NY, USA, 1992. ISBN 0-07-015305-1. xiv + 299 pp. LCCN QA76.76.R48 D36 1992. US$44.95.

German:1992:CLC

[47] Hallett German. Command language cookbook for mainframes, minicomputers, and PC’s: DOS/OS/2 batch language, Clist, DCL, Perl, and REXX. Van Nostrand Reinhold, New York, NY, USA, 1992. ISBN 0-442-00801-5. xiv + 352 pp. LCCN QA76.7 .G45 1992.

Goldberg:1992:RH

[48] Gabriel Goldberg and Philip H. Smith, III. The REXX handbook. J. Ranade IBM series. McGraw-Hill, New York, NY, USA, 1992. ISBN 0-07-023682-8. xxiv + 646 pp. LCCN QA76.73.R24 G65 1991.

IBM:1992:TEVa

[49] IBM Corporation, San Jose, CA, USA. TSO extensions version 2: procedures language MVS/REXX reference, sixth edition, October 1992. various pp. SC28-1883-5. File number: S370/S390-39.

IBM:1992:TEVb

[50] IBM Corporation, San Jose, CA, USA. TSO extensions version 2: procedures language MVS/REXX user’s guide, fourth edition, 1992. various pp. SC28-1882-3. File number: S370/390-39. Program number: 5685-025.

IBM:1992:TEVc

[51] IBM Corporation, San Jose, CA, USA. TSO extensions version 2: system diagnosis: procedures language MVS/REXX and CLIST, fourth edition, February 1992. various pp. LY28-1887-3. Restricted Materials of IBM. Program number: 5685-025.

Munro:1992:CRC

[52] Dick Munro. CMS and REXX course notes. Media Algonquin, Nepean, ON, Canada, 1992. ISBN 1-55076-165-X. 40 + 132 + 145 pp.

Anonymous:1993:VRP

[53] Anonymous. VX-Rexx for OS/2 (Programmer’s Guide and Reference) 2.0. Watcom International Corp., Waterloo, ON, Canada, 1993. ISBN 1-55094-074-0. 724 (est.) pp. LCCN ????

Anonymous:1993:WVRa

[54] Anonymous. WATCOM VX-REXX for OS/2 programmer’s guide and reference. Watcom International Corp., Waterloo, ON, Canada, 1993. ISBN 1-55094-068-6. xviii + 446 pp. LCCN ????

Anonymous:1993:WVRb

[55] Anonymous. WATCOM VX-REXX for OS/2 programmer’s guide and reference. Watcom International Corp., Waterloo, ON, Canada, version 2 edition, 1993. ISBN 1-55094-074-0. xxiv + 724 pp. LCCN QA76.73.R24.

Cepeda:1993:RBB

[56] Oscar Cepeda. OS/2 REXX from bark to byte. IBM redbooks. IBM International Technical Support Organization, Boca Raton, FL, USA, 1993. ISBN 0-7384-0572-8. ???? pp. LCCN QA76.73.R24.

Cowlishaw:1993:PR

[57] M. F. Cowlishaw. Personal REXX. Quercus Systems, 1993. Computer diskettes plus documentation.

Gargiulo:1993:RTE

[58] Gabriel F. Gargiulo. REXX in the TSO environment. The QED IBM mainframe series. QED Publishing Group, Boston, MA, USA, revised edition, 1993. ISBN 0-89435-418-3. xxi + 471 pp. LCCN QA76.73.R24 G37 1992. US$34.95.

Giguere:1993:ROW

[59] Eric Giguère. REXX and the OS/2 Workplace Shell. Dr. Dobb’s Journal of Software Tools, 18(13):44, 46, 48, December 1993. CODEN DDJOEB. ISSN 1044-789X.

Goran:1993:RRS

[60] Dick Goran. REXX Reference Summary Handbook (OS/2). CFS Nevada, Inc., CFS Nevada, Inc. 953 E. Sahara Avenue, Suite 9B Las Vegas, Nevada 89104-3012, USA, 1993. ISBN 0-9639854-0-X. 102 (est.) pp. LCCN ????

Hockware:1993:VRV

[61] Hockware Inc., ???? VisPro/REXX (Visual programming with REXX), 1993. 196 pp.

IBM:1993:ICLa

[62] IBM Corporation, San Jose, CA, USA. IBM compiler and library for SAA REXX/370: diagnosis guide, release 2.0, 1993. viii + 39 pp. SH19-8179-00.

IBM:1993:ICLb

[63] IBM Corporation, San Jose, CA, USA. IBM compiler and library for SAA REXX/370: user’s guide and reference, release 2.0, 1993. xvi + 197 pp. SH19-8160-01.

IBM:1993:TEVa

[64] IBM Corporation, San Jose, CA, USA. TSO extensions version 2 : REXX/MVS user’s guide, fifth edition, 1993. various pp. SC28-1882-04. File number: S370/390-39. Program number: 5685-025.

IBM:1993:TEVb

[65] IBM Corporation, San Jose, CA, USA. TSO extensions version 2: REXX/MVS reference, seventh edition, 1993. SC28-1883-06. File number: S370/390-39.

Kees:1993:RPG

[66] Peter Kees. REXX in der Praxis. (German [REXX in Practice]). R. Oldenbourg, München, Germany, 1993. ISBN 3-486-22666-5. 279 pp. LCCN ????

Kiesel:1993:RAT

[67] Peter C. Kiesel. REXX: advanced techniques for programmers. J. Ranade IBM series. McGraw-Hill, New York, NY, USA, 1993. ISBN 0-07-034600-3. xv + 239 pp. LCCN QA76.73.R24 K53 1993. US$39.95.

Nirmal:1993:RTT

[68] Barry K. Nirmal. REXX—tools and techniques. The QED IBM mainframe series. QED Pub. Group, Boston, MA, USA, 1993. ISBN 0-89435-417-5. xix + 264 pp. LCCN QA76.76.U84 N57 1992. US$39.95.

Cowlishaw:1994:EHR

[69] Mike Cowlishaw. The early history of REXX. IEEE Annals of the History of Computing, 16(4):15–24, Winter 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//dlib.computer.org/an/books/an1994/pdf/a4015.pdf;   http://
www.computer.org/annals/an1994/a4015abs.htm.

Gargiulo:1994:MR

[70] Gabriel F. Gargiulo. Mastering OS/2 REXX. Wiley, New York, NY, USA, 1994. ISBN 0-471-51901-4. xxi + 417 pp. LCCN QA76.73.R24 G37

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
;   http://www.loc.gov/catdir/bios/wiley047/94153212.html;
http://www.loc.gov/catdir/description/wiley032/94153212.html;
 http://www.loc.gov/catdir/toc/onix02/94153212.html.

Gargiulo:1994:UR

[71] Gabriel F. Gargiulo. Using OS/2 REXX. QED Pub. Group, Boston, MA, USA, 1994. ISBN 0-89435-449-3. various pp. LCCN QA76.73.R24.

German:1994:RHB

[72] Hallett German. OS/2 2.1 REXX Handbook: Basics, Applications, and Tips. VNR’s OS/2 series; VNR computer library. Van Nostrand Reinhold, New York, NY, USA, 1994. ISBN 0-442-01734-0 (paperback), 0-471-13132-6 (variant). xvii + 459 pp. LCCN QA76.73.R24 G47 1994.

Goran:1994:RRS

[73] Dick Goran. REXX Reference Summary Handbook (OS/2). CFS Nevada, Inc., CFS Nevada, Inc. 953 E. Sahara Avenue, Suite 9B Las Vegas, Nevada 89104-3012, USA, 1994. ISBN 0-9639854-1-8. 148 (est.) pp. LCCN ????

Kynning:1994:RPH

[74] Bengt Kynning. REXX Procedurspråk: hur du programmerar din PC med OS/2. (Swedish) [REXX Procedural Language: How you program your PC with OS/2]. Studentlitteratur, Lund, Sweden, 1994. ISBN 91-44-48541-7. ix + 300 pp. LCCN ????

Marks:1994:PLS

[75] Brian Marks. The programming language standards scene, ten years on: Paper 20: Rexx. Computer Standards and Interfaces, 16(5–6):555–557, 1994. CODEN CSTIEZ. ISSN 0920-5489.

Proffit:1994:MLC

[76] Brian Proffit. A multicolumn list-box container for OS/2. Dr. Dobb’s Journal of Software Tools, 19(5):90, 92–94, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Richardson:1994:WRP

[77] Ronny Richardson. Writing OS/2 REXX programs. IBM McGraw-Hill series. McGraw-Hill, New York, NY, USA, 1994. ISBN 0-07-052372-X (paperback). xxiii + 390 pp. LCCN QA76.73.R24 R53 1994. US$39.95.

Rudd:1994:ADU

[78] Anthony S. Rudd. Application development using OS/2 REXX. Wiley, New York, NY, USA, 1994. ISBN 0-471-60691-X (acid-free). xviii + 416 pp. LCCN QA76.76.A65 R83 1994. US$39.95.

Schindler:1994:TYR

[79] William F. Schindler and Esther Schindler. Teach yourself REXX in 21 days. SAMS Publishing, Indianapolis, IN, USA, 1994. ISBN 0-672-30529-

1. xx + 527 pp. LCCN QA76.73.R24 S35 1994. US$29.99 (US$39.99 Can.).

Sulger:1994:ERC

[80] Art Sulger. Extending REXX with C++. Dr. Dobb’s Journal of Software Tools, 19(10):36–37, 68, 86–88, September 1994. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1995:NPI

[81] Anonymous. New products: Interactive Unix Guide; SEDIT and S/REXX for Linux; Linux Internet Archives. Linux Journal, 16:??, August 1995. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827

(electronic). URL http://noframes.linuxjournal.com/lj-issues/
issue16/np16a.html;   http://noframes.linuxjournal.com/lj-issues/
issue16/np16b.html;   http://noframes.linuxjournal.com/lj-issues/
issue16/np16c.html.

Anonymous:1995:NPU

[82] Anonymous. New products: Unix Cockpit; InfoMagic and Lasermoon Release Moo-Tiff; CODEC for Linux; BGI Graphical Toolkit for Linux; Stock Charting Tool for Linux; Object REXX for Linux. Linux Journal, 15:??, July 1995. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827

(electronic). URL http://noframes.linuxjournal.com/lj-issues/
issue15/np15a.html;   http://noframes.linuxjournal.com/lj-issues/
issue15/np15b.html;   http://noframes.linuxjournal.com/lj-issues/
issue15/np15c.html;   http://noframes.linuxjournal.com/lj-issues/
issue15/np15d.html;   http://noframes.linuxjournal.com/lj-issues/
issue15/np15e.html;   http://noframes.linuxjournal.com/lj-issues/
issue15/np15f.html.

Callaway:1995:RCT

[83] Merrill Callaway. The REXX cookbook: a tutorial guide to the REXX language in OS/2 and Warp on the IBM personal computer. Whitestone, Albuquerque, NM, 1995. ISBN 0-9632773-4-0 (book); 0-9632773-5-9 (diskette). various pp. LCCN QA76.73.R24 C35 1995.

December:1995:HCU

[84] John December and Mark Ginsburg. HTML and CGI unleashed. SAMS Publishing, Indianapolis, IN, USA, 1995. ISBN 0-672-30745-6. xxvii + 830 pp. LCCN QA76.76.H94 D43 1995 Reserves. Contents: The World Wide Web as a communication system — A developer’s tour of the Web — Options for Web connections — Web development principles and methodology overview — Web planning — Web analysis – Web design — Web implementation — Web promotion – Web innovation — Design and implementation style and techniques — Basic HTML — Advanced HTML — Forms – Multimedia — Basics of imagemaps — Implementation tools — Development and language environments. Contents: Principles of gateway programming — Gateway programming fundamentals — Gateway programming I: programming libraries and databases — Gateway programming II: text search and retrieval tools – Scripting for the unknown: the control of chaos – Transaction security and security administration

– Gateway programming language options and a server modification case study — VRML on the Web — C-based gateway scripting — Writing CGI scripts in REXX — A Web coloring book — A campus-wide information system — A hypertext news interface — A graphical Web page counter. Contents: Appendixes: A. Sources of further information — B. HTML tag and supporting information summary — C. Glossary. PC and Macintosh compatible CD-ROM includes: HTML examples in the book, CGI scripts that you can use for forms, guestbooks, searching, and more, Web development tools such as HTML editors, graphics, and Internet search and retrieval applications.

Goran:1995:RRS

[85] Dick Goran. REXX Reference Summary Handbook. CFS Nevada, Inc., CFS Nevada, Inc. 953 E. Sahara Avenue, Suite 9B Las Vegas, Nevada 89104-3012, USA, third edition, 1995. ISBN 0-9639854-2-6. xiv + 208 pp. LCCN ????

Hallberg:1995:CH

[86] Bruce A. Hallberg and Kathy Ivens. OS/2 certification handbook. New Riders Publishing, Carmel, IN, USA, 1995. ISBN 1-56205-407-4. xxv + 1024 pp. LCCN QA76.76.O63 H3437 1995. REXX coverage by Bill Camarda.

IBM:1995:ITEa

[87] IBM Corporation, San Jose, CA, USA. IBM TSO extensions for MVS: REXX/MVS reference: version 2 release 5, eighth edition, 1995. various pp. C28-1883-07. File number: S370/390-39.

IBM:1995:ITEb

[88] IBM Corporation, San Jose, CA, USA. IBM TSO extensions for MVS: REXX/MVS user’s guide: version 2 release 5, sixth edition, December 1995. SC28-1882-05. File number: S370/390-39. Program number: 5685-025.

IX:1995:RSR

[89] iX Corporation, 575 W. Madison #3610, Chicago, IL 60661, USA. RXSQL — SQL for REXX, 1995. iii + 28 pp. URL http://www.wrkgrp.com/

support/manuals/rxsql.pdf.

Marco:1995:IRD

[90] Lou Marco. ISPF/REXX development for experienced programmers. CBM Books, Fort Washington, PA, USA, 1995. ISBN 1-878956-50-7. xvi + 201 pp. LCCN QA76.9.I58 M32 1995. US$35.00.

Richardson:1995:WVR

[91] Ronny Richardson. Writing VX-REXX programs. McGraw-Hill, New York, NY, USA, 1995. ISBN 0-07-911910-7, 0-07-911911-5 (paper). xxii + 336 pp. LCCN QA76.73.R24 R54 1995. US$55.00; US$39.95.

Yamada:1995:YRN

[92] Fumito Yamada. Yasashii Rexx Nyumon. (Japanese) [Introduction to Easy Rexx]. Daiwa Art, Tokyo, 1995. ISBN 4-88648-434-4. ???? pp. LCCN ????

ANSI:1996:ANSb

[93] American National Standards Institute and Information Technology Industry Council. American National Standard for information technology: programming language REXX: ANSI X3.274-1996. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1996. iv + 167 pp.

Cittadino:1996:RI

[94] Michael Cittadino. A REXX interpreter. Thesis (M.S.C.S.), North Central College, Napierville, IL, USA, 1996. ix + 275 pp.

Cowlishaw:1996:N

[95] Mike Cowlishaw. NetReXX. Web document, IBM Hursley, Hursley, UK,

1996. ???? pp. URL http://www.ibm.com/Technology/NetRexx/. Im-

plements a Rexx-language compiler into Java.

Rudd:1996:PUM

[96] Anthony S. Rudd. Practical usage of MVS REXX. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-19952-7 (Berlin : pbk.). xiv + 310 pp. LCCN QA76.73.R24 R82 1996.

Seki:1996:RJJ

[97] Takanori Seki. REXX Jiyu Jizai. (Japanese) [Handy with REXX]. Sciencesha, Tokyo, Japan, 1996. ISBN 4-7819-0810-1. ???? pp. LCCN ????

Veneskey:1996:ORE

[98] Gwen L. Veneskey, Will Trosky, and John J. Urbaniak. Object REXX by Example. Aviar Inc., Pittsburgh, PA, USA, 1996. ISBN 0-9652329-0-5. 313 + 2 + 10 pp. LCCN ???? Includes diskette.

Wittemann:1996:R

[99] Karlheinz Wittemann and Stefan Gerten. REXX unter OS/2. Hanser Fachbuch, München, Germany, 1996. ISBN 3-446-18508-9. xx + 575

pp. LCCN ???? URL http://www.bsz-bw.de/cgi-bin/ekz.cgi?
SWB04962238.

Yamashita:1996:RNY

[100] Naohiko Yamashita. REXX ni Yoru OS/2 Nyumon. (Japanese) [Introduction to OS/2 via REXX]. CQ Shuppan, Tokyo, Japan, 1996. ISBN 4-7898-3562-6. ???? pp. LCCN ????

comp-lang-rexx:1997:CLR

[101] comp.lang.rexx. Internet mailing list, 1997. URL news:comp.lang.rexx. The Usenet newsgroup comp.lang.rexx is in the general topic area of computer-related topics and covers the specific topic of the REXX command language. This group is an open forum. About 74% of the Usenet servers on the Internet subscribe to this group. The average number of messages sent to this group per day is 19 or about 33 kilobytes. About 4% of messages on this group also appear on other groups.

Cowlishaw:1997:NLa

[102] M. F. Cowlishaw. The NetRexx language. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-806332-X. viii + 197 pp. LCCN QA76.73.N47 C68 1997. See also supplement [118].

Cowlishaw:1997:NLb

[103] Mike Cowlishaw. The NetRexx language. IBM UK Laboratories, Hursley Park, Winchester, England, April 13, 1997. iv + 164 pp. URL http:/

/www-306.ibm.com/software/awdtools/netrexx/nrldef.pdf. Specifi-

cation 1.00.

Ender:1997:OOP

[104] Tom Ender. Object oriented programming with REXX. Wiley, New York, NY, USA, 1997. ISBN 0-471-11844-3 (paperback). xii + 253 pp. LCCN

QA76.64 .E52 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/;   http://www.loc.gov/catdir/toc/onix01/96045983.
html.

Goran:1997:RRS

[105] Dick Goran. REXX Reference Summary Handbook. CFS Nevada, Inc., CFS Nevada, Inc. 953 E. Sahara Avenue, Suite 9B Las Vegas, Nevada 89104-3012, USA, fourth edition, 1997. ISBN 0-9639854-3-4. 268 (est.) pp. LCCN ????

IBM:1997:CJA

[106] IBM. Creating Java Applications Using NetRexx. IBM Technical Support Organization, San Jose, CA, USA, September 1997. xx + 287 pp.

URL http://www.redbooks.ibm.com/pubs/pdfs/redbooks/sg242216.

pdf. SG24-2216-00.

Leibrock:1997:RLG

[107] Gerd Leibrock. Das REXX Lexikon: Begriffe, Anweisungen, Funktionen. (German) [The REXX Lexicon: Terms, Instructions, Functions]. R. Oldenbourg, München, Germany, 1997. ISBN 3-486-23672-5. 598 pp. LCCN ????

REXX-L:1997:R

[108] REXX-L. Internet mailing list, 1997. URL mailto:listserv@vmd.cso. uiuc.edu. REXX-L is a mailing list on the Rexx programming language. Membership in the group is available to anyone. The discussion archives are open to everyone. The organization responsible for running this group is the University of Illinois at Urbana-Champaign, USA. Their server name is VMD. CSO. UIUC. EDU. To join the group, send the message “sub REXX-L your name;; to listserv@vmd.cso.uiuc.edu.

Turton:1997:ORW

[109] Trevor Turton and Ueli Wahli. Object REXX for OS/2 Warp. IBM International Technical Support Organization, San Jose, CA, USA, 1997. ISBN 0-13-273467-2. xxx + 336 pp. LCCN QA76.64 .T87 1997.

VMREXX:1997:VR

[110] VM-REXX. Internet mailing list, 1997. URL mailto:listserv@vm. marist.edu. VM-REXX is a mailing list covering the topic of the VM/SP REXX programming language. Membership in the group is available to anyone. The discussion archives are open to everyone. The organization responsible for running this group is Marist College, Poughkeepsie, New York, USA. Their server name is VM. MARIST. EDU. To join the group, send the message “sub VM-REXX your name” to listserv@vm.marist.edu.

Wahli:1997:ORW

[111] Ueli Wahli, Ingo Holder, and Trevor Turton. Object REXX for Windows NT and Windows 95. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-858028-6. xxxv + 550 pp. LCCN QA76.64 .W34 1997.

Gargiulo:1998:RTC

[112] Gabriel F. Gargiulo. REXX—with OS/2, TSO & CMS features: quick reference guide. MVS Training Inc., Pittsburgh, PA, USA, 1998. ISBN 1-892559-03-X. vii + 96 pp. LCCN QA76.73.R24. URL http://www.

mvs-training.com.

Ousterhout:1998:SHL

[113] John K. Ousterhout. Scripting: Higher-level programming for the 21st century. Computer, 31(3):23–30, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Taylor:1998:IPU

[114] Pamela J. Taylor. Internet programming: Using NetRexx. Dr. Dobb’s Journal of Software Tools, 23(4):84–87, 101, April 1998. CODEN

DDJOEB. ISSN 1044-789X. URL http://www.ddj.com/ftp/1998/
1998_04/netrexx.txt.

Flatscher:1999:SIM

[115] Rony G. Flatscher. A syllabus for introducing MBA students to procedural and object-oriented programming (Object Rexx). In Haseman and Nazareth [134], page ?? ISBN ???? LCCN ????

Giguere:1999:WWR

[116] Eric Giguere, Rob Veitch, and J. Wesley Graham. WATCOM Windows REXX reference, Version ALPHA 2. WATCOM Publications, Waterloo, ON, Canada, 1999. v + 80 pp.

ANSI:2000:ANS

[117] American National Standards Institute and Information Technology Industry Council. American national standard for information technology: programming language REXX — amendments, errata and interpretations. 1430 Broadway, New York, NY 10018, USA, 2000. 14 pp. ANSI X3.274-1996/AM 1-2000. Approved December 13, 2000. Supplement to ANSI X3.274-1996.

Cowlishaw:2000:NLS

[118] Mike Cowlishaw. NetRexx Language Supplement. IBM UK Laboratories, Hursley Park, Winchester, England, August 23, 2000. iii +

45 pp. URL http://www-306.ibm.com/software/awdtools/netrexx/

nrlsupp.pdf. Version 2.00. This document is a supplement to [102].

Cowlishaw:2000:NUG

[119] Mike Cowlishaw. NetRexx User’s Guide. IBM UK Laboratories, Hursley Park, Winchester, England, August 31, 2000. ii + 40 pp. URL http://

www-306.ibm.com/software/awdtools/netrexx/nrusers.pdf. Version

2.00.

Prechelt:2000:ECS

[120] Lutz Prechelt. An empirical comparison of seven programming languages. Computer, 33(10):23–29, October 2000. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/books/co2000/pdf/rx023.pdf;   http://
www.computer.org/computer/co2000/rx023abs.htm.

Rudd:2000:PUT

[121] Anthony S. Rudd. Practical usage of TSO REXX. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2000. ISBN 1-85233-261-1. xv + 342 pp. LCCN QA76.73.R24 R84 2000.

Schindler:2000:ER

[122] William F. (Bill) Schindler. Down to Earth REXX. Perfect Niche Software, Scottsdale, AZ, USA, 2000. ISBN 0-9677590-0-5. xvii + 507 pp. LCCN ????

Cowlishaw:2003:NLO

[123] Mike Cowlishaw. NetRexx Language Overview. IBM UK Laboratories, Hursley Park, Winchester, England, February 24, 2003. ii +

21 pp. URL http://www-306.ibm.com/software/awdtools/netrexx/

nroverv.pdf. Version 2.01.

Prechelt:2003:SLG

[124] L. Prechelt. Are scripting languages any good? A validation of Perl, Python, Rexx, and Tcl against C, C++, and Java. Advances in Computers, 57:207–271, 2003. CODEN ???? ISSN 0065-2458.

Fosdick:2005:RPR

[125] Howard Fosdick. Rexx programmer’s reference. Wiley, New York, NY, USA, 2005. ISBN 0-7645-7996-7 (paperback), 0-7645-9695-0 (e-book). xxxiv + 686 pp. LCCN QA76.73.R24 F67 2005. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/;   http://media.wiley.com/
product_ancillary/67/07645799/DOWNLOAD/579967_appm.pdf; http:
//media.wiley.com/product_ancillary/67/07645799/DOWNLOAD/Rexx_
Prgrs_Reference_scripts.zip; http://www.loc.gov/catdir/toc/ecip055/
2004030560.html;   http://www.wrox.com/WileyCDA/WroxTitle/productCd-
0764579967,descCd-download_code.html.

Laird:2006:RER

[126] Cameron Laird and Kathryn Soraiz. Regular expressions: Rexx still going strong. UNIX review, ??(??):??, January 2006. CODEN UNRED5.

ISSN 0742-3136. URL http://www.unixreview.com/documents/s=
9953/ur0601h/ur0601h.html.

Dickey:1988:ACP

[127] Leroy J. Dickey and Lynne C. Shaw, editors. APL 88 Conference proceedings: APL, its past, its present, its future: 1–5 February 1988, Sydney, Australia, volume 18(2) of APL quote quad. ACM Press, New York, NY 10036, USA, 1988. ISBN 0-89791-253-5. ISSN 0163-6006. LCCN ???? ACM order number: 554880.

Dager:1990:PRS

[128] Cathie Dager and Nancy Larson, editors. Proceedings of the REXX Symposium for Developers and Users, June 11, 1990. Stanford Linear Accelerator Center (SLAC), Stanford, CA, USA, 1990. LCCN QA76.73.R49. SLAC report no. 368.

IEEE:1990:SPT

[129] IEEE, editor. Southeastcon ’90: proceedings: technologies today and tomorrow: April 1–4, 1990, Conference and Exhibit, Doubletree Hotel, New Orleans, Louisiana. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN ???? LCCN ???? Three volumes.

Dager:1991:PRS

[130] Cathie Dager, Dave Roos Gomberg, and Bebo White, editors. Proceedings of the REXX Symposium for Developers and Users: May 8–9, 1991, Asilomar, Pacific Grove, California. Stanford Linear Accelerator Center (SLAC), Stanford, CA, USA, 1991. LCCN QA76.73.R49. Second REXX Symposium. SLAC Report 379.

Anonymous:1992:PRS

[131] Anonymous, editor. Proceedings of the REXX Symposium for Developers and Users: May 3–5, 1992, Annapolis, Maryland. Stanford Linear Accelerator Center (SLAC), Stanford, CA, USA, 1992. LCCN QA76.73.R24. SLAC Report 401.

Anonymous:1993:PRS

[132] Anonymous, editor. Proceedings of the REXX Symposium for Developers and Users, May 18–20, 1993. Stanford Linear Accelerator Center (SLAC), Stanford, CA, USA, 1993. ISSN 1534-8954. LCCN ???? SLAC Report 422.

Anonymous:1994:PRS

[133] Anonymous, editor. Proceedings of the REXX Symposium for Developers and Users: May 1–3, 1994, Boston, Massachusetts. Stanford Linear Accelerator Center (SLAC), Stanford, CA, USA, 1994. LCCN QA76.73. SLAC report 447.

Haseman:1999:PFA

[134] W. David Haseman and Derek L. Nazareth, editors. Proceedings of the fifth Americas Conference on Information Systems: Milwaukee, Wisconsin REFERENCES 20

August 13–15, 1999. Association for Information Systems, Atlanta, GA, USA, 1999. ISBN ???? LCCN ????